OpenStudy (vera_ewing):

Write a conditional statement from the statement. A horse has 4 legs. a. If an animal has 4 legs then it is a horse. c. If an animal is a horse then it has 4 legs. b. Every horse has 4 legs. d. An animal has 4 legs and the animal is a horse. @satellite73

OpenStudy (vera_ewing):

@satellite73 I think it is B...

OpenStudy (anonymous):

conditional looks like if something then something

OpenStudy (vera_ewing):

Ohhh so it is A! IF, THEN!

OpenStudy (anonymous):

i would go with C if an animal is a horse, then it has four legs

OpenStudy (vera_ewing):

Ok, C...but why not A?

OpenStudy (anonymous):

because the statement "if an animal has four legs, then it is a horse" is silly my cat has four legs

OpenStudy (anonymous):

C

OpenStudy (vera_ewing):

Oh, ok. Thanks! Can you help with one more?

OpenStudy (anonymous):

if i can

OpenStudy (vera_ewing):

M is the midpoint of line AN, A has coordinates (5, 7), and M has coordinates (–1, 2). Find the coordinates of N.

OpenStudy (anonymous):

\[\frac{5+x}{2}=-1\] and also \[\frac{7+y}{2}=2\] solve for \(x\) and \(y\)

OpenStudy (vera_ewing):

ok i don't know what the first step would be...

OpenStudy (anonymous):

multiply by 2 to get rid of the fractions i will do the first you can do the second

OpenStudy (vera_ewing):

ok i am doing the second one right now

OpenStudy (anonymous):

\[\frac{5+x}{2}=-1\\ 5+x=-2\\ x=-7\]

OpenStudy (vera_ewing):

y=-3

OpenStudy (vera_ewing):

Thanks for the help! :)